For Problems 9 and 10: In a three-factor experiment suppose that factor A has 3 levels, factor B has 2 levels, and factor C has 3 levels. Also, for each of the combination of levels of the factors, 3 observations were measured. The researcher fitted the full model with main effects, all two-factor interactions, and all three-factor interactions.

9. How many possible treatments are there?

A. 8

B. 11

C. 18

D. 54

10. If all treatments were applied, what are the degrees of freedom (df) for Error in the ANOVA table?

A. 16

B. 22

C. 36

D. 108
Answer #1

(9)

number of treatment possible =levels of factor A*levels of factor B*levels of factor C

=3*2*3

=18

(10)

df (error )=levels of factor A*levels of factor B*levels of factor C*(observation-1)

=3*2*3*(3-1)

=18*2

=36

------------------------------------------------------------------------

DEAR STUDENT,

IF YOU HAVE ANY QUERY ASK ME IN THE COMMENT BOX,I AM HERE TO HELPS YOU.PLEASE GIVE ME POSITIVE RATINGS

*****************THANK YOU*****************